OpenStudy (anonymous):

Really need the answer, please help! The polar coordinates (8, 7π/4) represent the Cartesian point a. (4, 4) b. (4√2, 4√2) c. (-4√2, 4√2) d. (4√2, -4√2)

OpenStudy (phi):

what is 7 pi/4 radians translated to degrees ?

OpenStudy (anonymous):

401.1?

OpenStudy (phi):

multiply by 180/pi \[ \frac{7 \pi}{4} \cdot \frac{180}{\pi} \] notice that the pi s will cancel. multiply top times top and bottom times bottom simplify.

OpenStudy (anonymous):

315

OpenStudy (phi):

yes 315º next, draw x and y axes  if you start at the x-axis and go counter-clockwise (up and to the left) 315 degrees, where do you end up ? remember 360º is a full circle. straight up is 90º, then sideways is 180º, straight down is 270º...

OpenStudy (anonymous):

in the 4th quadrant

OpenStudy (phi):

Here is an example going 270º